OpenStudy (anonymous):

Which of these has the greatest area? if you can, could you use 3.14 for pi? rhombus with side lengths: 5 inches, height: 6 inches triangle with base length: ten inches, height: sixteen inches circle with diameter: 5 inches trapezoid with base lengths: 9 inches and eleven inches, height: 6 inches

OpenStudy (nurali):

Area of rhombus = L*H = 5*6 = 30 Area of triangle = 1/2 B*H = 10*16*1/2 = 80 Area of Circle = pi*r^2 = pi*5^2 = pi*25 = 78.5 Area of trapezoid = H * (L1+L2)/2 = 6 *(9+11)/2 = 60 So the triangle has the greatest area.

OpenStudy (anonymous):

ooooh ok, thank you so much :3